Circular No. 7951

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002ei IN MCG −01-09-024 M. Papenkova and W. D. Li, University of California, Berkeley (UCB),

report the LOTOSS (cf. IAUC 7906) discovery of an apparent supernova on unfiltered CCD images taken with KAIT on Aug. 8.5 (mag ∼ 16.5) and 9.4 UT (mag ∼ 16.6). SN 2002ei is located at α = 3^{h}16^{m}00^{s}.33, δ = −5^{o}30′07′′.4 (equinox 2000.0), which is 3′′.6 west and 10′′.1 south of the nucleus of MCG −01-09-024. A KAIT image of the same field on Feb. 2.2 UT showed nothing at this position (limiting mag ∼ 19.0).

Filippenko, R. J. Foley (UCB), Papenkova and B. Swift (UCB) report that inspection of CCD spectra (range 310–1000 nm) obtained on Aug. 9 UT with the Shane 3-m reflector at Lick Observatory reveals that the object is a supernova of type II in its plateau phase. Hα has a well-developed P-Cyg profile; many Fe II and other lines with such profiles are also visible.

SUPERNOVA 2002ej IN NGC 6324 T. Puckett, Mountain Town, GA; and B. Kerns, Stanton, ND, report

the discovery of an apparent supernova (mag 17.5) on an unfiltered CCD frame (limiting mag 19.3) taken with the Puckett Observatory 0.35-m automated supernova patrol telescope on Aug. 9.11 UT. The candidate is located at α = 17^{h}05^{m}22^{s}.62, δ = +75^{o}24′24′′.2 (equinox 2000.0), which is 9′′.8 west and 1′′.3 south of the center of NGC 6324. The new object was also present and confirmed on unfiltered CCD frames taken on Aug. 10.05, but it does not appear on unfiltered images taken by Puckett on July 16 and 24 (limiting mag ∼ 20.5) or on Palomar Sky Survey images taken on 1955 May 20 (limiting mag ∼ 20.0), 1992 July 28 and 1993 Aug. 19 (limiting mag ∼ 21.0).

COMETS C/2002 H8, C/2002 K8, C/2002 K9 AND C/2002 K10 (SOHO)

Further to IAUC 7948, D. Hammer reports measurements for comets detected on C2 SOHO website images by R. Kracht (C/2002 H8, Meyer group), R. Matson (C/2002 K8 and K9, Kreutz group) and D. Evans (C/2002 K10, Kreutz group). The reduced astrometry and orbital elements by the undersigned appear on the cited MPEC s.

Comet  2002 UT  α_{2000}  δ_{2000}  MPEC
C/2002 H8 Apr. 20.811  1^{h}58.^{m}5  +12^{o}54′  2002-P41
C/2002 K8 May 25.643  4 11.2  +19 27  2002-P17
C/2002 K9  25.976  4 12.1  +19 30  2002-P17
C/2002 K10  30.976  4 31.6  +20 11  2002-P17
2002 August 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Brian G. Marsden